### v2

@@ Abstract
Recent studies have investigated whether large language models (LLMs) can support obscure communication that requires specialized skills, such as inferring subtext or doublespeak. To conduct the investigation, researchers have used social deduction games (SDGs) as their experimental environment, in which players conceal and infer specific information. However, prior work has often overlooked how LLMs should be evaluated in such settings. Specifically, we point out two issues with the evaluation methods they employed. First, metrics used in prior studies are coarse-grained as they are based on overall game outcomes that often fail to capture event-level behaviors; Second, error analyses have lacked structured methodologies capable of producing insights that meaningfully support evaluation outcomes. To address these issues, we propose a macroscopic and systematic approach to the investigation. Specifically, we introduce seven fine-grained metrics that resolve the first issue. To tackle the second issue, we conducted a thematic analysis and identified four major reasoning failures that undermine LLMs’ performance in obscured communication.
@@ I Introduction
Large language models (LLMs) have shown that they can successfully conduct typical informative communication with humans by leveraging their natural language understanding abilities [1, 2, 3]. Inspired by these reports, there is a growing interest in adapting LLMs for obscured communication as a next step [4]. Here, the obscured communication indicates that the intended meanings of messages present in the communication are obscured or made unclear to understand, usually by controlling the spread of information. For example, in the real world, an imaginary role-play requires obscure communication to boost immersion in such a role-play. Among the required skills, doublespeak is one of the representative skills that enable a participant to achieve an advantageous position in such a play. So, since LLMs with doublespeak ability can support the play, researchers have recently focused on investigating such abilities. [5]
@@
Meanwhile, investigating whether LLMs can conduct obscure communication requires a specialized environment. As ordinary LLM use cases focus on providing information, obscured communication cannot be tested with such ordinary cases. To set a more suitable environment for the examination, researchers have used social deduction games (SDGs) because SDGs force players to use doublespeak to achieve their goals. Specifically, they let LLMs play an SDG and collected the game logs, including LLMs’ decisions and conversations. After, they devised several metrics using particular game actions for evaluating LLMs’ ability of doublespeak. [6] Also, they conducted qualitative analyses to discover erroneous actions or utterances. For example, [7] examined whether LLMs can support obscured communication by letting them play a famous SDG, Werewolf. However, to thoroughly examine LLMs’ doublespeak abilities, these attempts have two issues: (1) Macroscopic quantitative evaluation and (2) Non-systematic qualitative analysis.
@@
First, the issue of macroscopic evaluation refers to the tendency of prior works to use broad game-level outcomes as their metrics instead of specific event-level actions. Existing research often relies on simple metrics, such as win rates, which fail to account for the events that occur during gameplay [6, 7]. While these studies acknowledge the importance of certain skills for supporting obscured communication, they may overlook the need to measure these skills separately with such granular metrics. To scrutinize a specific skill, we suggest using microscopic analysis augmented with several fine-grained metrics. Specifically, we propose seven metrics to evaluate two essential skills regarding doublespeak: (1) Subtext inference, the deductive ability to identify doublespeak of others, and (2) Deceptive control, which is the ability to control information to make others confused.
@@
Second, the non-systematic evaluation issue indicates that previous studies lacked structured methods for error analysis. While certain studies identified abnormal patterns through simple qualitative analysis, the reproducibility of their results was limited by their analytical procedures. For instance, some studies employed LLMs to analyze data without incorporating human-annotation procedures [8, 9], while others did not adequately discuss their qualitative analysis methods for defining abnormal patterns. Also, they simply reported the occurrence tendency of the patterns rather than discussing how they affected their evaluation results [10]. We believe that conducting qualitative analysis in a systematic way provides both reproducibility and interoperability of experimental results. Therefore, we adopt a systematic approach known as thematic analysis [11] to identify recurring reasoning errors that hinder LLMs’ doublespeak. By correlating these qualitative patterns with our quantitative metrics, we uncover how specific reasoning failures impact LLMs’ performance in subtext inference and deceptive control.
@@
We conducted two studies to address these issues. In Study 1 (Section IV), we perform a quantitative analysis using our fine-grained metrics. The result reveals how coarse-grained metrics have overlooked finer aspects of doublespeak. In Study 2 (Section V), we conduct a thematic analysis on collected logs. Based on the result, we deduce several categories and link these findings to the results of Study 1. In both studies, we let four LLMs play SpyGame, a variant of simple SDG called SpyFall (Section III-A). Consequently, this study makes the following contributions:
  • We demonstrate the effectiveness of using microscopic quantitative analysis to assess the doublespeak of LLMs. Specifically, using seven fine-grained metrics, we show possibility of studying doublespeak in LLMs’ SDG play.
  • We identify four categories of abnormal reasoning patterns that harm obscured communication. Also, by combining results from the two studies, we discover a correlation between the results of the two analyses.
@@ II Related Works
This section reviews the existing literature in two ways. First, we outline how recent research has shifted its focus from evaluating LLMs on informative communication tasks to more complex settings that require obscured communication. Then, we point out two issues posed in their assessment: macroscopic evaluation and non-systematic analysis.
@@ II-A Adapting LLMs for Obscured Communication
LLMs have shown human-level performance in various informative communications. For example, researchers widely tested LLMs’ ability to question-answering or summarization [12, 13]. To accurately and reliably provide answers, LLMs for informative communication attempt to leverage true knowledge and suppress false statements during the generation process [14, 15]. So, researchers attempt to build an honest and harmless assistant for informative communication.
@@ II-B Attempts to assess LLMs in obscured communication
However, we suspect that such metrics are too coarse-grained to adequately capture intermediate behaviors during gameplay. Specifically, the win rate is only available post hoc, limiting its utility for formative assessment. Additionally, the action success rate may fail to account for the multi-player dynamics of SDGs in which performance involves not only task execution but also managing others’ perceptions, such as avoiding suspicion. In other words, achieving a higher win rate or action success rate in an SDG does not entail that the model possesses superior competence across all sub-skills involved in obscured communication.
Other researchers proposed more granular metrics, such as logicality or social behaviors, that can be computed by scoring the reasoning or statements made by LLMs during an Avalon game[9, 8]. However, such metrics proposed in those studies rely on an auxiliary LLM to qualitatively assess reasonings or statements, rather than being grounded in objective in-game events. Therefore, fine-grained quantitative metrics grounded in objective in-game events are essential for evaluating LLMs’ abilities in multi-agent obscured communication, such as subtext inference and deceptive control.
@@
In summary, the existing literature has limitations in its quantitative and qualitative analysis methods. To discover whether LLMs can support obscured communication, quantitative analysis with fine-grained metrics and validation of qualitative analysis are necessary.
@@ III-A1 Team formation
Similar to Spyfall, SpyGame requires two teams: six citizens and one spy. The citizens share a common location, called a hidden location, with each citizen playing a specific character in the location. For example, let us assume that the location is a school. Then, the citizens are assigned to distinct characters, such as a student or a principal. Meanwhile, the spy does not have any information about the location or characters.
@@ III-A2 Game Flow
From the second turn, the spy player guesses the hidden location by herself (Figure 1c). For each turn, she tells her guess to the system secretly and provides a number indicating how certain she is. When the certainty is high enough [^2] In our implementation, we use 9 as the threshold out of 10., the system announces her guess and ends the game. Otherwise, the game continues.
@@ III-B Models and Data
We selected four LLMs for our experiment: GPT-4[1], GPT-3.5-turbo (ChatGPT), Gemini-pro 1.0[3], and Llama2-70b-chat[2]. As obscured communication requires reasoning on subtexts, we selected large enough models with sufficient reasoning ability on simple tasks. Specifically, we selected models with a parameter size over 70 billion and a reasoning performance exceeding 70% accuracy[21] in 2023. LLMs provide their actions or utterances with their reasoning in natural language. So, we employed GPT-4 to sanitize their outputs into a machine-readable format.
To collect data, we let each tested model play SpyGame as a spy against strong or weak citizens. Here, the strong or weak citizen models were fixed based on performance reported in prior studies [21, 22, 1]: GPT-4 for strong and Llama2-70b-chat for weak citizens. And, we conducted our experiment repeatedly. Specifically, we used seven locations borrowing from the original game. For each location, we ran 24 games4442 types of citizens $\times$ 4 spy LLMs $\times$ 3 trials.. Consequently, we collected a total of 168 gameplay logs. Table I lists statistics of collected logs. On average, each game took 5.4 turns, with 23.8 reasoning steps which refers to a unit of LLM’s reasoning for taking an action.
TABLE I: Average statistics for all players in the collected logs
@@
IV Microscopic Quantitative Analysis
Our study aimed to investigate whether LLMs can support obscured communication by conducting a microscopic analysis of their SDG playing. We designed seven metrics that enable fine-grained analysis of two key SDG facets: subtext inference and deceptive control. Compared with most metrics adopted by previous research, which were designed on game-level outcomes, our approach allows for an analysis of intermediate behaviors during the game. Also, our quantitative metrics are fully automated metrics that do not require any manual labeling. In this section, we first discuss whether existing simple metrics are sufficient for measuring intermediate behaviors of LLMs. Then, we demonstrate our analysis can discover new insights that could not be captured by using current metrics by discussing our quantitative results.
IV-A Insufficiency of Previous Analysis
Previous studies commonly adopted macroscopic analyses to examine SDG gameplay behavior[8, 20, 19, 7, 6]. The most frequently used metrics are the win rates (WR) and living rounds (LR). These metrics evaluate a player’s performance based on final game outcomes: WR is determined by the win/loss result, while LR reflects how long the player remained active during the game.
@@ IV-A1 Metric result
Tables II (a) and III (a) present the results of the coarse-grained metrics when playing against strong and weak citizens, respectively. The WR and LR exhibited different trends in SDG play performance. In terms of WR, GPT-4 showed the highest performance against strong citizens (80.95%), followed by Gemini (57.14%), Llama (52.38%), and GPT-3.5 (33.33%). A similar order of performance was observed for against weak citizens. However, in terms of LR, Llama demonstrated the highest performance against strong citizens (9.00), followed by Gemini (6.95), GPT-4 (5.80), and GPT-3.5 (3.71). This trend is similar when examining the performance playing against weak citizens. Note that these metrics yield different results regarding the best-performing model: WR indicates GPT-4 as the best, while LR identifies Llama.
This discrepancy is due to multiple winning conditions specified in the SDG. These conditions allow players to win games in several ways. Therefore, to accurately identify the skills that contribute to winning, it is necessary to consider multiple abilities simultaneously. However, since WR and LR consider only the final state of the game, these metrics cannot examine different abilities separately. For example, in SpyGame, LR indicates how well a spy survives during the game. However, a longer survival does not always guarantee victory, leading to potential discrepancies between LR and WR. Hence, a microscopic method is required for a comprehensive analysis of LLM’s performance in various facets.
TABLE II: Quantitative results for playing against strong citizens
(a) General (b) Subtext inference (c) Deceptive Control
Winning Living Guess Notice Information Information Caught Vote Vote
Model Rate (%) Round (#) Success (%) Rate (%) Catching Deduction (%) Rate (%) Rate (%) Entropy
Gemini 57.14 6.95 50.00 $3$ 30.00 022.22 35.7 0.348
(14) (20) (9) (42)
GPT-3.5 33.33 3.71 27.78 28.57 - 28.57 007.14 25.0 0.730
(18) (21) (14) (12)
GPT-4 80.95 5.80 81.25 61.90 - 61.90 025.00 33.3 0.953
(16) (21) (4) (30)
Llama2 52.38 9.00 - 09.52 - 00.00 100.00 45.2 0.634
(0) (19) (10) (126)
Total 55.95 6.365 35.78 $3$ 30.86 037.83 40.47 0.642
Metrics whose lower values indicate better results.
Note that Information Catching was not measured playing against strong citizens.
TABLE III: Quantitative results for playing against weak citizens
(a) General (b) Subtext inference (c) Deceptive control
Winning Living Guess Notice Information Information Caught Vote Vote
Model Rate (%) Round (#) Success (%) Rate (%) Catching (%) Deduction Rate (%) Rate (%) Entropy
Gemini 90.48 5.28 85.71 090.48 61.11 - 00.00 07.1 0.420
(14) (18) (2) (42)
GPT-3.5 71.43 2.52 71.43 076.19 92.86 - $0$
(21) (14) (6) (0)
GPT-4 90.48 2.42 90.48 100.00 87.50 - $0$
(21) (16) (2) (0)
Llama2 76.19 7.85 60.00 071.43 20.00 - 60.00 13.5 0.605
(5) (15) (5) (96)
Total 82.14 4.51 80.32 084.52 65.07 - 20.00 11.59 0.548
Metrics whose lower values indicate better results.
Note that Information Deduction was not measured playing against weak citizens.
IV-B Subtext Inference

### v3

@@ Abstract
Recent studies have investigated whether large language models (LLMs) can support obscured communication, which is characterized by core aspects such as inferring subtext and evading suspicions. To conduct the investigation, researchers have used social deduction games (SDGs) as their experimental environment, in which players conceal and infer specific information. However, prior work has often overlooked how LLMs should be evaluated in such settings. Specifically, we point out two limitations with the evaluation methods they employed. First, metrics used in prior studies are coarse-grained as they are based on overall game outcomes that often fail to capture event-level behaviors; Second, error analyses have lacked structured methodologies capable of producing insights that meaningfully support evaluation outcomes. To address these limitations, we propose a microscopic and systematic approach to the investigation. Specifically, we introduce six fine-grained metrics that resolve the first issue. To tackle the second issue, we conducted a thematic analysis and identified four major reasoning failures that undermine LLMs’ performance in obscured communication.
@@ I Introduction
Large language models (LLMs) have shown that they can successfully conduct typical informative communication with humans by leveraging their natural language understanding abilities [1, 2, 3]. Inspired by these reports, there is a growing interest in adapting LLMs for obscured communication as a next step [4]. Here, the obscured communication indicates that the intended meanings of messages present in the communication are obscured or made unclear to understand, usually by controlling the spread of information. For example, in the real world, an imaginary role-play requires obscure communication to boost immersion in such a role-play. Among the required skills, doublespeak is one of the representative skills that enable a participant to achieve an advantageous position in such a play. So, since LLMs with doublespeak ability can support the play, researchers have recently focused on investigating such abilities [5].
@@
In this context, evaluating LLMs’ ability to support obscured communication is challenging because it requires assessing not just informativeness or task completion, but also the ability to strategically conceal sensitive information and infer hidden meanings from indirect cues. Unlike cooperative dialogue, where participants work toward shared goals, obscured communication is inherently adversarial, making it a significantly more complex and nuanced form of interaction. To set a more suitable environment for the examination, researchers have used social deduction games (SDGs) because SDGs force players to use doublespeak to achieve their goals. Specifically, they let LLMs play an SDG and collected the game logs, including LLMs’ decisions and conversations. After, they devised several metrics using particular game actions for evaluating LLMs’ ability of doublespeak [6]. Also, they conducted qualitative analyses to discover erroneous actions or utterances. For example, [7] examined whether LLMs can support obscured communication by letting them play a famous SDG, Werewolf. However, we question whether these prior attempts are sufficient to thoroughly assess whether LLMs can effectively support obscured communication due to their two limitations: (1) Macroscopic quantitative evaluation and (2) Non-systematic qualitative analysis.
@@
First, the issue of macroscopic evaluation refers to the tendency of prior works to use broad game-level outcomes as their metrics, rather than specific event-level actions. Existing research often relies on simple metrics, such as win rate, which fail to account for the events that occur during gameplay [6, 7]. While these studies acknowledge the importance of specific aspects for supporting obscured communication, they may overlook the need to measure these aspects separately with such granular metrics. This raises our first research question (RQ1): Does a high score based on coarse-grained metrics imply that an LLM supports all the key aspects? To address this question, we propose six metrics to evaluate two core aspects that characterize obscured communication: (1) Subtext inference, the deductive ability to identify doublespeak of others, and (2) Deceptive control, the ability to manage one’s information to mislead others through effective doublespeak.
@@
Second, the non-systematic evaluation issue indicates that previous studies lacked structured methods for error analysis. While certain studies identified abnormal patterns through simple qualitative analysis, the reproducibility of their results was limited by their analytical procedures. For instance, some studies employed LLMs to analyze data without incorporating human-annotation procedures [8, 9], while others did not adequately discuss their qualitative analysis methods for defining abnormal patterns. Also, they simply reported the occurrence tendency of the patterns rather than discussing how they affected their evaluation results [10]. We believe that conducting qualitative analysis in a systematic way provides both reproducibility and interoperability of experimental results. In this context, we raise our second research question (RQ2): What types of reasoning errors can be systematically identified through qualitative analysis, and how do these errors hinder LLMs from supporting obscured communication? To address this question, we employ a systematic approach known as thematic analysis [11] to identify recurring reasoning errors that hinder LLMs’ doublespeak. By correlating these qualitative patterns with our quantitative metrics, we uncover how specific reasoning failures impact LLMs’ performance in subtext inference and deceptive control.
@@
In Study 1 (Section IV), we perform a quantitative analysis using our fine-grained metrics. The result reveals how coarse-grained metrics have overlooked finer aspects of doublespeak. In Study 2 (Section V), we conduct a thematic analysis on collected logs. Based on the result, we deduce several categories and link these findings to the results of Study 1. In both studies, we let four LLMs play SpyGame, a variant of simple SDG called SpyFall (Section III-A). Consequently, this study makes the following contributions:
  • We propose a fine-grained evaluation framework for analyzing LLMs’ doublespeak in obscured communication. Specifically, our proposed metrics enable the assessment of detailed skills that cannot be evaluated using coarse-grained metrics.
  • We identify four categories of reasoning failures that hinder LLMs from supporting obscured communication, and show that these categories provide explanatory support for our quantitative results.
@@ II Related Works
This section reviews the existing literature in two ways. First, we outline how recent research has shifted its focus from evaluating LLMs on informative communication tasks to more complex settings that require obscured communication. Then, we point out two limitations posed in their assessment: macroscopic evaluation and non-systematic analysis.
@@ II-A Adapting LLMs for Obscured Communication
LLMs have shown human-level performance in various informative communications. For example, researchers widely tested LLMs’ ability in question-answering or summarization [12, 13]. To accurately and reliably provide answers, LLMs for informative communication attempt to leverage true knowledge and suppress false statements during the generation process [14, 15]. So, researchers attempt to build an honest and harmless assistant for informative communication.
@@ II-B Attempts to assess LLMs in obscured communication
However, we suspect that such metrics are too coarse-grained to capture intermediate behaviors during gameplay adequately. Specifically, the win rate is only available post hoc, limiting its utility for formative assessment. Additionally, the action success rate may fail to account for the multi-player dynamics of SDGs in which performance involves not only task execution but also managing others’ perceptions, such as avoiding suspicion. In other words, achieving a higher win rate or action success rate in an SDG does not entail that the model possesses superior competence across all skills involved in core aspects of obscured communication.
Other researchers proposed more granular metrics, such as logicality or social behaviors, that can be computed by scoring the reasoning or statements made by LLMs during an Avalon game[9, 8]. However, such metrics proposed in those studies rely on an auxiliary LLM to qualitatively assess reasoning or statements, rather than being grounded in objective in-game events. Therefore, fine-grained quantitative metrics grounded in objective in-game events are essential for evaluating LLMs’ abilities in multi-agent obscured communication, such as subtext inference and deceptive control.
@@ III Experiment
TABLE I: Overview of players in SpyGame
# Player Objectives Required Abilities
Spy 1 Identify the hidden location while avoiding detection Concealing or Revealing Information
Citizen 6 Detect the spy without revealing the location
@@ III-A1 Team formation
As illustrated in Table I, SpyGame requires two teams: six citizens and one spy. The citizens share a common location, called a hidden location, with each citizen playing a specific character in the location. For example, let us assume that the location is a school. Then, the citizens are assigned to distinct characters, such as a student or a principal. Meanwhile, the spy does not have any information about the location or characters.
@@ III-A2 Game Flow
From the second turn, the spy player guesses the hidden location by herself (Figure 1c). For each turn, she secretly tells the system her guess and provides a number indicating her level of certainty. When the certainty is high enough [^2] In our implementation, we use 9 as the threshold out of 10., the system announces her guess and ends the game. Otherwise, the game continues.
@@ III-B Models and Data
We selected four LLMs for our experiment: GPT-4[1], GPT-3.5-turbo (ChatGPT), Gemini-pro 1.0[3], and Llama2-70b-chat[2]. As obscured communication requires reasoning on subtexts, we selected models that are large enough and have sufficient reasoning ability on simple tasks. Specifically, we selected models with a parameter size over 70 billion and a reasoning performance exceeding 70% accuracy[21] in 2023. LLMs provide their actions or utterances with their reasoning in natural language. So, we employed GPT-4 to sanitize their outputs into a machine-readable format.
To collect data, we let each tested model play SpyGame as a spy against strong or weak citizens. Here, the strong or weak citizen models were fixed based on performance reported in prior studies [21, 22, 1]: GPT-4 for strong and Llama2-70b-chat for weak citizens. And, we conducted our experiment repeatedly. Specifically, we used seven locations borrowing from the original game. For each location, we ran 24 games4442 types of citizens $\times$ 4 spy LLMs $\times$ 3 trials.. Consequently, we collected a total of 168 gameplay logs. Table II lists statistics of collected logs. On average, each game took 5.4 turns, with 23.8 reasoning steps, which refers to a unit of LLM’s reasoning for taking an action.
TABLE II: Average statistics for all players in the collected logs
@@
IV Quantitative Analysis with Fine-Grained Metrics
TABLE III: Overview of tested metrics
Metric Precondition Level Measuring Point
Prior Win Rate - Coarse-grained Outcome of game
Metrics Living Round - Coarse-grained How long survived
Guess Success Rate Spy win Mid-grained How subtext inference ability contributed to win
Information Catching Rate Exposure occured Fine-grained How successfully exploited explicit clue
Our Information Deduction Rate Exposure not occured Fine-grained How successfully exploited implicit clue
Metrics Caught Rate Spy lose Mid-grained How deceptive control ability contributed to lose
Vote Rate Day/Final vote occured Fine-grained How successfully avoided concentrated suspicion
Vote Entropy Day/Final vote occured Fine-grained How successfully dispersed suspicion
To investigate RQ1 regarding the adequacy of coarse-grained metrics in capturing LLMs’ abilities, we conduct a fine-grained analysis of model behavior in SDG. To this end, we design six quantitative metrics, which are presented in Table III, targeting two core aspects of obscured communication subtext inference and deceptive control.
For each aspect, we propose one mid-grained metric to assess its overall contribution to the spy’s success or failure, and two fine-grained metrics to evaluate how well the model performs the associated skills. From the following subsection, we first discuss the limitations of existing simple metrics, which focus solely on game-level outcomes and are insufficient for capturing the intermediate behaviors exhibited by LLMs during gameplay. We then demonstrate how our proposed metrics uncover insights into the way LLMs utilize specific reasoning skills in obscured communication, based on our quantitative analysis.
Additionally, we performed statistical analyses to assess the degree of model-level differences, with the expectation that such variation may offer additional interpretive insights. To examine differences between models within each metric, we conducted statistical analyses. Given the small size of 168 games, we used non-parametric statistical tests. For metrics composed of continuous variables, we performed the Kruskal-Wallis test[23] to examine group differences. For metrics consisting of categorical variables, we used the chi-squared test[24] to assess differences between groups.
IV-A Limitations of Existing Metrics
Previous studies commonly adopted macroscopic analyses to examine SDG gameplay behavior[8, 20, 19, 7, 6]. The most frequently used metrics are the win rate (WR) and living round (LR). These metrics evaluate a player’s performance based on final game outcomes: WR is determined by the win/loss result, while LR reflects how long the player remained active during the game.
@@ IV-A Limitations of Existing Metrics
    $WR=\frac{\#\text{ victory}}{\#\text{ total games}}$ (1)
@@
    $LR=\frac{\sum\#\text{ turns spy survived}}{\#\text{ total games}}$ (2)
@@ IV-A1 Metric result
Tables V (a) and VI (a) present the results of the coarse-grained metrics when LLMs played against strong and weak citizens, respectively. When playing against strong citizens, statistical differences across models were found in both WR ($\chi^{2}=9.81$, $p<0.05$) and LR ($H=42.09$, $p<0.001$) metrics: GPT-4 achieved the highest WR (80.95%), followed by Gemini (57.14%), Llama (52.38%), and GPT-3.5 (33.33%). However, when playing against weak citizens, no significant difference was found in WR across models ($\chi^{2}=4.14$, $p=0.24$). We suspect that weak citizens’ low gameplay proficiency made it easier for all models to achieve high win rates as the spy. Unlike WR, LR revealed a different trend. Llama consistently recorded the longest survival time (9.00 and 7.85), even outperforming GPT-4, which had the highest WR. This discrepancy between WR and LR illustrates that success in surviving longer does not necessarily translate to overall victory. It suggests that each metric captures different aspects of gameplay, potentially tied to distinct skills
Furthermore, this discrepancy highlights that coarse-grained metrics are challenging to interpret because both WR and LR are influenced by a range of overlapping factors, including guessing accuracy, voting dynamics, and the amount of information citizens disclose. As a result, it remains unclear which specific underlying skills these metrics are actually measuring. Hence, a microscopic method is required for a comprehensive analysis of LLM’s performance in various facets.
IV-B Analysis of Subtext Inference
We designed three metrics to enable a microscopic analysis of the first core aspect, subtext inference. This aspect concerns the model’s ability to uncover hidden information, either through deduction or by capturing clues embedded in the conversation. In particular, subtext inference has two skills: information capture and information deduction.